OpenStudy (anonymous):

Is it this correct? 3+8/9 + 2+2/7 = 56/63 and then I added and got 5+74/63

hartnn (hartnn):

3+2=5 now \[5+\frac{8}{9}+\frac{2}{7}=\frac{63*5+8*7+2*9}{63}\] got this?

OpenStudy (anonymous):

how did you get the stuff after =

OpenStudy (anonymous):

\[3+\frac{8}{9}+2+\frac{2}{7}==5+\left(\frac{8}{9}+\frac{2}{7}\right)=5+\left(\frac{74}{63}\right) \]

OpenStudy (anonymous):

Yes I got that. But it isn't the answer.

OpenStudy (anonymous):

The correct answer is 6 and 11/63

OpenStudy (anonymous):

\[5+\left(\frac{74}{63}\right)=5+\left(1+\frac{11}{63}\right)=6+\frac{11}{63} \]